OpenStudy (noahbred):

Please help fasttt i will medal and fan Simplify: (–8) • (6) – (5) • (–2) A. 58 B. 16 C. –38 D. –48

OpenStudy (♪chibiterasu):

We follow PEMDAS. You know what that is right?

OpenStudy (noahbred):

wait wrong equation sorryyyy

OpenStudy (noahbred):

Simplify: –4(x – 2) – 5(y + 2) + (–3)(8 – 6b) A. –4x – 5y + 18b + 6 B. –4x – 5y + 18b – 6 C. –4x – 5y – 18b – 22 D. –4x – 5y + 18b – 26

OpenStudy (noahbred):

@confluxepic

OpenStudy (♪chibiterasu):

Break them down. –4(x – 2) – 5(y + 2) (–3)(8 – 6b)

OpenStudy (noahbred):

ohh its A yeaaa

OpenStudy (noahbred):

@pooja195 help?

pooja195 (pooja195):

(−8)(6)−(5)(−2) −48−(5)(−2) −48−(−10)=?

OpenStudy (noahbred):

-58?

pooja195 (pooja195):

-48 + 10 = ? -+-= +

OpenStudy (noahbred):

-58

OpenStudy (♪chibiterasu):

Adding a positive to a negative will reduce the number

pooja195 (pooja195):

No we are ADDING 10 to -48

pooja195 (pooja195):

48 - 10 =?

OpenStudy (noahbred):

-38

pooja195 (pooja195):

well done :D

OpenStudy (noahbred):

so its C